OpenStudy (anonymous):

What are the exact solutions of x2 − 5x − 1 = 0? x = the quantity of 5 plus or minus the square root of 29 all over 2 x = the quantity of negative 5 plus or minus the square root of 29 all over 2 x = the quantity of 5 plus or minus 2 times the square root of 5 all over 2 x = the quantity of negative 5 plus or minus 2 times the square root of 5 all over 2

OpenStudy (anonymous):

x = the quantity of 5 plus or minus the square root of 29 all over 2

OpenStudy (anonymous):

Using the quadratic formula: \[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[x = \frac{-(-5) \pm \sqrt{(-5)^2-4(1)(-1)}}{2 \times 1}\] \[x = \frac{5 \pm \sqrt{25+4}}{2}\] \[x = \frac{5 \pm \sqrt{29}}{2}\]